A jar contains 36 coins all quarters and nickels. The value of the coins is $5.10. How many of each coin is in the jar
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52799)

.
A jar contains 36 coins all quarters and nickels. The value of the coins is $5.10. How many of each coin is in the jar
~~~~~~~~~~~~~~~~~~~~~~~~~~

This problem does not have a solution.

Such a combination of 36 quarters and nickels that is worth $5.10 doesn't exist.

So, there are only two options:

a)  EITHER the problem doesn't have a solution,
b)  OR the input data is wrong.

Now the question is: for what reason do you liter the forum with wrong problems?


Answer by Alan3354(69443)

A jar contains 36 coins all quarters and nickels. The value of the coins is $5.10. How many of each coin is in the jar
-----
q + n = 36
25q + 5n = 510 --> 5q + n = 102
==============
5q + n = 102
q + n = 36
------------------ Subtract
4q = 66
q = 16.5 quarters
n = 19.5 nickels
-------
I'm not sure that 1/2 of a quarter is worth 12.5 cents.